# Long Grounded Thoughts: Distilling Compositional Visual Reasoning Chains at Scale

### v1

@@ Abstract
Recent progress in multimodal reasoning has been driven largely by undisclosed datasets and proprietary data synthesis recipes, leaving open questions about how to systematically build large-scale, vision-centric reasoning datasets, particularly for tasks that go beyond visual math. In this work, we introduce a new reasoning data generation framework spanning diverse skills and levels of complexity with over 1M high-quality synthetic vision-centric questions. The dataset also includes preference data and instruction prompts supporting both offline and online RL. Our synthesis framework proceeds in two stages: (1) scale, where imagery and metadata (captions, bounding boxes) are used to generate diverse, verifiable visual questions; and (2) complexity, where a composition hardening algorithm merges simpler questions from the previous stage into harder, still verifiable visual problems. Reasoning traces are synthesized through a two-stage process that leverages VLMs and reasoning LLMs, producing CoT traces for VLMs that capture the richness and diverse cognitive behaviors found in frontier reasoning models. Remarkably, we show that finetuning Qwen2.5-VL-7B on our data outperforms all open-data baselines across all evaluated vision-centric benchmarks, and even surpasses strong closed-data models such as MiMo-VL-7B-RL on V*Bench, CV-Bench and MMStar-V. Perhaps most surprising, despite being entirely vision-centric, our data transfers positively to text-only reasoning (MMLU-Pro, +2.98%) and audio reasoning (MMAU, +1.32%), demonstrating its effectiveness. Similarly, despite no containing videos or embodied visual data, we observe notable gains (+10%) when evaluating on a single-evidence embodied QA benchmark (NiEH). Finally, we use our data to analyze the entire VLM post-training pipeline. Our empirical analysis highlights that (i) SFT on high-quality data with non-linear reasoning traces is essential for effective online RL, (ii) staged offline RL matches online RL’s performance while reducing compute demands, and, (iii) careful SFT on high quality data can substantially improve out-of-domain, cross-modality transfer.
Figure 1: Overview of our two-stage synthesis framework. First, we synthesize multiple-choice questions (MCQs) from dense captions and grounded object metadata, emphasizing scale and diversity while teaching basic cognitive behaviors (verification, backtracking, correction). Later, we harden questions by composing them into visual reasoning problems that requires decomposition and higher-order reasoning. For each stage, we also synthesize reasoning traces by first distilling CoTs from VLMs and then expanding them with reasoning LLMs, yielding traces that are in the distribution of VLM outputs yet richer in reasoning depth.
@@ 1 Introduction
Since the arrival of DeepSeek R1 (deepseekai2025deepseekr1incentivizingreasoningcapability), a wealth of open-source reasoning datasets has been developed for language-based reasoning (muennighoff2025s1simpletesttimescaling; jung2025prismatic; openthoughts; lin2025goedelproverv2scalingformaltheorem), as innovations in data curation and distillation have proven to be powerful levers for advancing open-source models—sometimes even surpassing closed ones in specific domains such as mathematics. In contrast, open-source multimodal reasoning efforts lag behind, perhaps because it is not entirely obvious how to apply and synthesize long chain-of-thoughts (CoTs) with complex reasoning structures (e.g., verification, backtracking, subgoal setting) for vision-centric reasoning tasks. Indeed, most open vision-centric reasoning datasets with complex reasoning structures remain either limited in scale or focused on visual math. Table 1 highlights this by comparing popular multimodal reasoning datasets that exhibit traces with complex structures.
A recent step forward came from LongPerceptualThought (LPT; liao2025longperceptualthoughts), which synthesized a 30K dataset of long structured reasoning traces for vision-centric tasks. By combining VLMs with reasoning LLMs, their work produced traces with complex reasoning structures.
@@
In this work, we push this line of research further with a synthesis framework that tackles three core challenges: scale, complexity and richness of the reasoning trace, and we introduce Long Grounded Thoughts. Long Grounded Thoughts is one of the largest vision-centric reasoning datasets to date, distilled at scale using frontier reasoning VLMs and LLMs. Our data contains over 1M SFT examples spanning multiple levels of complexity and reasoning traces distilled from models such as Qwen2.5-VL-72B, Qwen3-235B-A22B-Thinking, and R1-671B. We also present 130K examples for offline and online RL, supporting the full spectrum of post-training.
@@
We show that finetuning a 7B VLM on our data outperforms all open-data baselines on several challenging vision-centric benchmarks. Remarkably, our model outperforms strong closed-data models like MiMO-VL-7B-RL in 3 out of 5 benchmarks, and even surpasses proprietary systems such as GPT-4o and Claude 3.7 on a few vision-centric benchmarks. Notably, even without any video or embodied-QA data, our fine-tuned model achieves substantial improvement (+10 points vs base model) on a modified NiEH single-evidence QA task (kim2025needlesembodiedhaystackenvironment). Perhaps more surprising, despite being entirely vision-centric, our data also transfers positively to text-only reasoning (MMLU-Pro) and audio reasoning (sound and music) on an Omni-7B model.
@@
Finally, we use our data to systematically analyze post-training in VLMs, revealing the following findings: (i) online RL necessitates prior “skill teaching”—instruct models without cognitive skills underperform SFT on high-quality synthetic data; (ii) multi-staged offline training (SFT→DPO) achieves nearly the same performance as online RL with less compute and higher scalability; (iii) online RL yields early gains but plateaus around 70K examples; (iv) surprisingly, high quality SFT alone can substantially improve out-of-domain, cross-modality transfer.
@@
Table 1: Comparison of our visual reasoning dataset with prominent open-source counterparts. Our dataset scales to over 1M+ examples. Cognitive behaviours are marked as present () if the average count on a sampled subset of 1000 examples is at least 10%, we quantify the behaviour following the methodology from gandhi2025cognitivebehaviorsenableselfimproving; liao2025longperceptualthoughts.
@@
Cognitive Behaviours Dataset # QA Primary Domain Data Type Subgoal Backtrack Verify PixMo-AskModelAnything (deitke2024molmo) 162K Visual Question Answer SFT SCI-Reason (ma2025scireason) 12.6k Scientific Image Reasoning SFT LENS (wang2024multi) 40k Multi-Scenario Visual Reasoning SFT DriveLMM-o1 (ishaq2025drivelmm)  22k Driving Visual Reasoning SFT Virgo (du2025virgo)  19K Visual Math SFT VLLA-Thinking (chen2025sft)  152K Multimodal Reasoning & Visual Math SFT/RL LongPerceptualThoughts (liao2025longperceptualthoughts) 30k Vision-Centric Reasoning SFT/RL Ours 1M+ Vision-Centric Reasoning & Compositional Reasoning SFT/RL
@@ 2 Method
Our data generation framework, illustrated in Figure 1, tackles three core challenges: scale, complexity, and richness of the reasoning trace. In the first stage, we generate large-scale MCQs using high-quality captions and grounded metadata (i.e., bounding boxes). This stage emphasizes scale and diversity, while teaching models basic cognitive behaviors such as verification, backtracking, and self-correction (Section 2.1). In the second stage (Section 2.2), we apply a composition hardening algorithm that merges MCQs from the first stage into more challenging, multi-hop problems that require decomposition and higher-order reasoning, where the basic skills act as building blocks. For each stage, we then synthesize reasoning traces (Section 2.3). For the synthesis, in the first stage, we focus on scale and rely on models around the 30B parameter scale (e.g., Qwen2.5-VL-7B, R1-32B). In the second stage, where problems demand decomposition and chaining of skills, we leverage the most powerful open VLMs and LLMs (e.g., Qwen2.5-VL-72B, R1-671B). In both cases, we distill first CoTs from VLMs and subsequently expand them with richer trajectories using reasoning LLMs. This allows to obtain reasoning traces that are in the VLM distribution but contain the richness of the reasoning model. In our notation, $\mathcal{M}_{\text{VLM}}$ denotes the VLM, $\mathcal{M}_{\text{LLM}}$ the LLM, and $\mathcal{M}_{\text{Reason}}$ the reasoning LLM. Importantly, we assume access to a pool of natural images with highly descriptive captions that include fine-grained visual details such as DOCCI (OnoeDocci2024).
@@ 2.1 Scale and Diversity
Figure 2: Scaling Behaviour of LPT vs Ours for SFT. We find that using additional metadata (here bounding boxes) in addition to highly details captions allows for more diverse and controlled generation of MCQ successfully scaling beyond 1M+ examples.
In this stage, we focus on synthesizing visual problems (i.e., multiple-choice questions, MCQs) at scale. Two key requirements for scaling MCQ synthesis are simplicity and diversity. Simplicity ensures that we can generate a large number of MCQs in a cost-efficient manner, while diversity guarantees that each question is unique and non-redundant. A straightforward approach is to provide an LLM with a highly detailed description of the image and prompt it to generate MCQs based on the image and its associated dense descriptions. This setup offers two advantages for later stages: (1) it ensures that each question is answerable using only the dense descriptions, allowing us to synthesize reasoning traces at a later stage purely in the text modality, and (2) the multiple-choice format enables explicit verification of correctness, which is essential for constructing positive and negative preference pairs. This strategy follows prior work such as LongPerceptualThought (liao2025longperceptualthoughts).
However, when scaling this approach from thousands to millions of datapoints, we found that question diversity quickly saturates (see Figure 2). We hypothesize that this occurs because at scale the LLM collapses and repeatedly generates similar MCQs about the same objects, when conditioned only on global captions.
To strengthen visual grounding and increase diversity, we incorporate additional object-level metadata (e.g., bounding boxes and class tags) to guide MCQ generation. This encourages the LLM to formulate questions that are about a specific object of interest. In our scenario, we assume object metadata is not given. Thus each image is first processed with the Grounded-Segment-Anything model (ren2024grounded), producing open-vocabulary bounding boxes and object tags. We then augment the global caption with the coordinates and tag of a selected object. This combined information is passed to the generator model using a structured prompt(see suppl material), which forces the model to generate a question that targets the specified region of the image. Formally, given an image v with dense textual descriptions c, and object-level metadata omd (bounding-box coordinates and object tags), our goal is to construct a triplet:
@@
where q is the generated question and a is the correct answer. Intuitively, for an image with $K$ detected objects, we can produce $K$ distinct MCQs by conditioning on each object. Interestingly, we found that including normalized bounding-box coordinates, in addition to tags, further improved question grounding, even though the LLM itself operates purely in the text modality. Finally, we apply a rigorous filtering protocol to ensure quality. Each generated MCQ further undergoes automated checks to validate its format and correct answer. We refer to this as Stage 1 (A.1 for details).
Figure 3: Reasoning trace comparison between our model (post-SFT and RL) and the vanilla base model. Both models initially fail to identify the dog in the image. The base model terminates with an incorrect answer based on this flawed premise. In contrast, our model demonstrates a non-linear reasoning process; it employs self-verification and backtracking to challenge and self-correct its initial assessment. This correction appears to stem from a trace where the model relies on captioning and grounding as a bridge between language and vision; notably grounding on the dog triggers the revised path on a second "self-captioned" verification structure. This behavior is notable as captions were not explicitly included in the training traces, perhaps suggesting captioning and grounding as part of the thinking process could be an emergent capability of training on our data.
2.2 Complexity via Composition
A limitation of synthetic MCQ generation using detailed captions and object metadata is that the resulting questions are often relatively easy. In practice, we observed a large percentage of these problems can be solved directly by the base VLMs. As shown in Figure 4(a), our metadata driven synthesis strategy already produces questions that are harder than LPT-style caption-only generation, yet many remain solvable for the base model. Interestingly, we observe that VLM performance improves even on these simpler problems, likely because the augmented reasoning traces reinforce basic cognitive skills such as self-verification and backtracking.
@@ 2.2 Complexity via Composition
where $\textnormal{q}$$^{\star}$ is the synthesized complex question and $\textnormal{a}$$^{\star}$ its corresponding solution. Ideally, for a VLM model being able to solve the new synthesized problem, it should break it down in the set of simpler questions (see the suppl material for prompts and examples). We refer to this data as Stage 2.
(a) Complexity estimation via multiple rollouts.
(b) Analysis of Cognitive Behaviors in CoTs.
Figure 4: Analysis of our data splits. (a) Complexity estimation via multiple rollouts on synthesized MCQs using Qwen2.5-VL as a policy. Darker green color represents easier problems. (b) Analysis of Cognitive Behaviors in CoTs. Our data exhibits higher frequencies of subgoal setting, backtracking, and verification, indicating a more deliberate and structured reasoning process. Estimation of cognitive behaviours and terminology was borrowed from  gandhi2025cognitivebehaviorsenableselfimproving.
@@ 2.3 Synthesizing Reasoning Traces
To generate long CoTs familiar to the VLM, we follow the strategy presented in LPT (liu2025prorlprolongedreinforcementlearning). Specifically, we prompt a VLM $\mathcal{M}_{\text{VLM}}$ with the image and its corresponding MCQ to produce a rationale (${\textnormal{z}}_{1}$) and final prediction (${\textnormal{a}}_{1}$), denoted as $({\textnormal{z}}_{1},{\textnormal{a}}_{1}):=\mathcal{M}_{\text{VLM}}({\textnormal{v}},{\textnormal{q}})$. Sampling from a VLM ensures that the synthesized CoTs remain within distribution, as in LPT (liao2025longperceptualthoughts), which we find crucial for downstream performance. Specifically, we find naively sampling from $\mathcal{M}_{\text{Reason}}$ often produces CoTs that deviate significantly from those of the VLM, which we experimentally observed significantly degrades downstream performance during fine-tuning. Thus, we adopt the thought-expansion mechanism from LPT, guiding $\mathcal{M}_{\text{Reason}}$ to extend rather than rewrite the VLM’s rationale. Formally, we structure the prompt as:
$\text{User: }{\textnormal{c}}\oplus{\textnormal{q}},\$$\text{Assistant: }\texttt{<think>}\oplus{\textnormal{z}}_{1}$
and ask $\mathcal{M}_{\text{Reason}}$ to continue the thought, producing $({\textnormal{z}}_{2},{\textnormal{a}}_{2})$, here $\oplus$ denotes concatenation. This preconditioning allows the reasoning LLM to expand familiar traces while injecting richer, non-linear problem-solving strategies. Figure 4(b) analyzes cognitive behaviors in our reasoning traces, highlighting notable non-linear patterns in our data, especially in hardened MCQs from stage 2.
Scaling challenge. At larger scales, we observed that $\mathcal{M}_{\text{Reason}}$ frequently referenced the provided image description explicitly (e.g., “the image description says,…”), rather than producing an expansion as if grounded in the visual input. LPT handled this with aggressive filtering. At scale, filtering ratios quickly became a bottleneck, reducing throughput. To overcome this, we introduce a guided decoding mechanism on the sampling algorithm that uses regular expressions (regex) to constrain generations. This substantially improves distillation efficiency and utilization at scale.
Multi-Stage Synthesis. For MCQs generated in Stage 1 (simpler problems), we use Qwen2.5-VL-7B for the initial VLM synthesis and DeepSeek-R1-Distill-Qwen-32B as our $\mathcal{M}_{\text{Reason}}$ for expansion. For Stage 2 (harder, composed problems), we reserve the strongest available models—Qwen2.5-VL-72B, R1-671B and Qwen3-235B-Thinking for reasoning LLM expansion.
3 Offline and Online Synthesis for RL
To build a preference dataset for offline RL, we follow (setlur2024rlincorrectsyntheticdata; zhang2025backtracking; kimiteam2025kimik15scalingreinforcement) and define pairwise comparisons based on correctness and compactness. Formally, we define:

### v2

# Long Grounded Thoughts: Synthesizing Visual Problems and Reasoning Chains at Scale

@@ Abstract
Despite rapid progress, multimodal reasoning still lacks a systematic approach to synthesize large-scale vision-centric datasets beyond visual math. We introduce a framework able to synthesize vision-centric problems spanning diverse levels of complexity, and the resulting dataset with over 1M high-quality problems including: reasoning traces, preference data, and instruction prompts supporting SFT, offline and online RL. Our vision-centric synthesis framework uses a two-stage process focusing on: (1) generating diverse verifiable questions from existing images at scale, and (2) creating complex compositional visual problems by merging simpler questions. Remarkably, finetuning Qwen2.5-VL-7B on our data outperforms existing open-data baselines across evaluated vision-centric benchmarks, and our best configurations match or surpass strong closed-data models such as MiMo-VL-7B-RL on V*Bench, CV-Bench and MMStar-V. Notably, despite being entirely vision-centric, our data transfers positively to text-only reasoning (MMLU-Pro, +3.7%) and audio reasoning (MMAU, +1.32%), demonstrating its effectiveness. Similarly, despite containing no embodied visual data, we observe notable gains (NiEH, +8.8%) when evaluating open-ended embodied QA. Lastly, we use our data to comprehensively analyze at scale (1M+) the entire VLM post-training pipeline showing that (i) SFT on high-quality data with cognitive behaviors on reasoning traces is essential to scale online RL, (ii) offline RL could match online RL’s performance while disaggregating compute demands, and, (iii) SFT on high quality data also improve out-of-domain, cross-modality transfer.
Machine Learning, ICML
Figure 1: Overview of our two-stage data generation framework. First, we synthesize multiple-choice questions (MCQs) from dense captions and grounded object metadata, emphasizing scale and diversity. Later, we harden questions by composing them into visual reasoning problems that requires decomposition and higher-order cognitive patterns. For each stage, we also synthesize reasoning traces by first distilling CoTs from VLMs and then expanding them with reasoning LLMs, yielding traces that are in the distribution of VLM outputs yet richer in reasoning depth. The caption is never directly seen by the trained model.
@@ 1 Introduction
Since the arrival of DeepSeek R1 (DeepSeek-AI et al., 2025), a wealth of open-source reasoning datasets has been developed for language-based reasoning (Muennighoff et al., 2025; Jung et al., 2025; Team, 2025; Lin et al., 2025), as innovations in data curation and distillation have proven to be powerful levers for advancing open-source models. In contrast, open-source multimodal reasoning efforts seem to lag behind. Perhaps because it is not entirely obvious how to synthesize new vision-centric problems with an appropriate level of complexity and long chain-of-thoughts (CoTs) with complex reasoning structures (e.g., verification, backtracking, subgoal setting) specifically for vision-centric reasoning tasks. Indeed, most open vision-centric reasoning datasets with complex problems and reasoning structures remain either limited in scale or focused on visual math. Table 1 highlights this by comparing popular multimodal reasoning datasets that exhibit traces with complex structures.
A recent step forward came from LongPerceptualThought (LPT; Liao et al., 2025a), which synthesized 30K problems with structured reasoning traces for vision-centric tasks. By using dense captions as a proxy between images and language, LPT combines VLMs with reasoning LLMs to synthesize new problems and traces with complex cognitive structures. However, we identify two critical bottlenecks in their framework when pushed to scale: 1) problem synthesis saturation, relying solely on captions for problem synthesis limits the diversity and grounding of the generated vision-centric questions, causing downstream performance to saturate quickly and 2) cognitive simplicity, the synthesized reasoning traces for those problems—while richer than previous works—still lack higher-order structures.
@@
In this work, we push this line of research further with a synthesis framework that tackles these three core challenges: scale, complexity, and CoT richness. We introduce Long Grounded Thoughts, containing 1M+ new vision-centric problems with complex reasoning traces distilled from frontier models (e.g., Qwen2.5-VL, R1-671B) including 129K additional problems for offline and online RL.
Compared to recent methods like LPT (Liao et al., 2025a), our framework delivers significant improvements:
  • Scale: We overcome the bottlenecks of caption only synthesis methods such as LPT by leveraging grounded object metadata. This allows us to scale problem synthesis from 30K to over 1M+ high-quality vision centric problems without performance saturation.
[…]
@@
We show that finetuning a 7B VLM on our data outperforms all open-data baselines on several vision-centric benchmarks. Notably, our best configurations match or outperform strong closed-data models like MiMO-VL-7B-RL in 3 out of 5 benchmarks. Even without any video or embodied-QA data in our dataset, our fine-tuned model achieves substantial improvement on an open-ended embodied QA task (Kim and Ammanabrolu, 2025). Perhaps more surprising, despite being entirely vision-centric, our data also transfers positively to text-only reasoning (MMLU-Pro) and audio reasoning (sound and music) on an Omni-7B model. We additionally show that training on our verifiable multiple choice problems also generalize to open ended visual tasks.
@@
Finally, we leverage our dataset to conduct a comprehensive analysis of the full VLM post-training spectrum for vision-centric tasks. To the best of our knowledge, this represents one of the largest studies of its kind at scale on VLMs to date revealing four critical insights: (i) online RL necessitates offline “teaching” of cognitive behaviours—instruct models without cognitive skills underperform SFT on high-quality synthetic data; (ii) multi-staged offline training (SFT→DPO) could reach competitive performance compared to online RL with decoupled compute and higher scalability; (iii) online RL on base models yields early gains but plateaus quickly; (iv) high quality SFT alone can substantially improve out-of-domain, cross-modality transfer.
@@
Table 1: Comparison of our visual reasoning dataset with prominent open-source counterparts. Our dataset scales to over 1M+ examples. Cognitive behaviours are marked as present () if the average count on a sampled subset of 1000 examples is at least 10%, we quantify the behaviour following the methodology from Gandhi et al. (2025); Liao et al. (2025a).
@@
Cognitive Behaviours Dataset # QA Primary Domain Data Type Subgoal Backtrack Verify PixMo-AskModelAnything (Deitke et al., 2024) 162K Visual Question Answer SFT SCI-Reason (Ma et al., 2025) 12.6k Scientific Image Reasoning SFT LENS (Wang, 2024) 40k Multi-Scenario Visual Reasoning SFT DriveLMM-o1 (Ishaq et al., 2025)  22k Driving Visual Reasoning SFT Virgo (Du et al., 2025)  19K Visual Math SFT VLLA-Thinking (Chen et al., 2025b)  152K Multimodal Reasoning & Visual Math SFT/RL LongPerceptualThoughts (Liao et al., 2025a) 30k Vision-Centric Reasoning SFT/RL Ours 1M+ Vision-Centric Reasoning & Compositional Reasoning SFT/RL
@@ 2 Method
Our data generation framework, illustrated in Figure 1, tackles three core challenges: scale, complexity, and richness of the reasoning trace. In the first stage, we generate large-scale multiple choice visual problems (MCQs) using high-quality captions and grounded metadata (i.e., bounding boxes). This stage emphasizes synthesis of diverse problem at scale (Section 2.1). In the second stage (Section 2.2), we apply a composition hardening algorithm that merges MCQs from the first stage into more challenging, multi-hop problems that require decomposition and higher-order reasoning. For each stage, we then synthesize reasoning traces that could be used for SFT and offline RL (Section 2.3). In our notation, $\mathcal{M}_{\text{VLM}}$ denotes the VLM, $\mathcal{M}_{\text{LLM}}$ the LLM, and $\mathcal{M}_{\text{Reason}}$ the reasoning LLM.
@@ 2.1 Scale and Diversity
We focus on synthesizing visual problems (i.e., multiple-choice questions, MCQs) at scale. Two key requirements for scaling MCQ synthesis are simplicity and diversity. Simplicity ensures that we can generate a large number of MCQs in a cost-efficient manner, while diversity guarantees that each question is unique and non-redundant. A straightforward approach is to provide an LLM with a highly detailed description of the image and prompt it to generate MCQs based on the image and its associated dense descriptions.
[…]
Figure 2: Scaling Behaviour of LPT vs Ours for SFT. We find that using additional metadata (here bounding boxes) in addition to highly details captions allows for more diverse and controlled generation of MCQs, successfully scaling beyond 1M+ examples. Please refer to appendix A.2 for additional analysis on diversity.
Formally, given an image v with dense textual descriptions c, and object-level metadata omd (bounding-box coordinates and object tags), our goal is to construct a triplet:
@@
where q is the generated question and a is the correct answer. Intuitively, for an image with $K$ detected objects, we can produce $K$ distinct MCQs by conditioning on each object. Interestingly, we found that including normalized bounding-box coordinates further improved question grounding, even though the LLM itself operates purely in text . Finally, we apply filtering protocol to ensure quality. See Stage 1 (A.1 for details). We assume object metadata is not given and first processed images with Grounded-SAM (Ren et al., 2024), producing open-vocabulary bounding boxes and object tags.
(a) Complexity estimation via multiple rollouts.
(b) Analysis of Cognitive Behaviors in CoTs.
Figure 3: Analysis of our data splits. (a) Complexity estimation via multiple rollouts on synthesized MCQs using Qwen2.5-VL as a policy. Darker green color represents easier problems. (b) Analysis of Cognitive Behaviors in CoTs. Our data exhibits higher frequencies of subgoal setting, backtracking, and verification, indicating a more deliberate and structured reasoning process. Estimation of cognitive behaviours and terminology from  (Gandhi et al., 2025). Table 13 shows detailed cognitive behaviours analysis on several datasets.
Figure 4: Reasoning trace comparison between our model (post-SFT and RL) and the vanilla base model. Both models initially fail to identify the dog in the image. The base model terminates with an incorrect answer based on this flawed premise. In contrast, the model finetuned on Long Grounded Thoughts demonstrates a non-linear reasoning process; it employs self-verification and backtracking to challenge and self-correct its initial assessment. This correction appears to stem by relying on captioning and grounding as a bridge between language and vision; notably grounding on the dog triggers the revised path on a second verification structure.
2.2 Compositionality for Complexity
A limitation of synthetic MCQ generation is that the resulting questions are often relatively easy. In practice, we observed a large percentage of these problems can be solved directly by the base VLMs. As shown in Figure 3(a) and Table 3, our simple grounded synthesis strategy (Stage 1) already generates problems that are harder than caption-only methods, yet many remain solvable for the base model.
@@ 2.2 Compositionality for Complexity
where $\textnormal{q}$$^{\star}$ is the synthesized complex question and $\textnormal{a}$$^{\star}$ its corresponding solution. (see A.4 for prompts and examples). We refer to this data as Stage 2. We define multi-step reasoning as a reasoning trace that decomposes the original problem into manageable sub-steps to reach the final answer (i.e. subgoal setting). As shown in Figure 3(b) stage 2 leads to a significantly higher subgoal setting. Figure 3(a), on the other hand shows the generated questions are notably harder.
@@ 2.3 Synthesizing Reasoning Traces
We synthesize long CoTs, using the two stage expansion strategy presented in LPT (Liao et al., 2025a). We prompt a VLM $\mathcal{M}_{\text{VLM}}$ with the image and its corresponding MCQ to produce a rationale (${\textnormal{z}}_{1}$) and final prediction (${\textnormal{a}}_{1}$), denoted as $({\textnormal{z}}_{1},{\textnormal{a}}_{1}):=\mathcal{M}_{\text{VLM}}({\textnormal{v}},{\textnormal{q}})$. Then, we structure a prompt as:
$\text{User: }{\textnormal{c}}\oplus{\textnormal{q}},\text{Assistant: }\texttt{<think>}\oplus{\textnormal{z}}_{1}$
and ask $\mathcal{M}_{\text{Reason}}$ to continue the thought, producing $({\textnormal{z}}_{2},{\textnormal{a}}_{2})$, here $\oplus$ denotes concatenation. Figure 3(b) analyzes cognitive behaviors in our reasoning traces, highlighting notable non-linear patterns in our data, especially in hardened MCQs from stage 2. We emphasize the caption is never seen for the training model or during inference. The caption is used only by the reasoning LLM during the expansion as the LLM cannot "see".
Do we need expanded long CoTs? Yes. We ablated the long CoT expansion strategy on 750K examples (Table 12). Surprisingly, training on simplified ‘Short CoT‘ distilled only from VLM traces degrades performance below even ‘No CoT‘ (direct answers), likely due to negative transfer from low-fidelity reasoning. Only ‘Long CoT’ traces obtained by sampling from the VLM and expanded with complex cognitive behaviors via the reasoning LLM yields consistent gains, validating the necessity of the expansion.
Scaling challenge. At larger scales, we observed that $\mathcal{M}_{\text{Reason}}$ frequently referenced the provided image description explicitly (e.g., “the image description says,…”), rather than producing an expansion as if grounded in the visual input. To overcome this, we introduce a guided decoding mechanism on the sampling algorithm that uses regular expressions (regex) to constrain generations. This substantially improves distillation efficiency and utilization at scale.
Multi-Stage Synthesis. For MCQs generated in Stage 1 (simpler problems), we use Qwen2.5-VL-7B for the initial VLM synthesis and DeepSeek-R1-Distill-Qwen-32B as our $\mathcal{M}_{\text{Reason}}$ for expansion. For Stage 2 (harder, composed problems), we reserve the strongest available models—Qwen2.5-VL-72B, R1-671B and Qwen3-235B-Thinking for reasoning LLM expansion. Crucially, 75% of our data uses accessible models (Stage 1) yet yields strong performance (Table 11). We release both subsets separately to enable both reproducibility and high-quality research without dependence on frontier teachers.
2.4 Offline and Online Synthesis for RL
To build a preference dataset for offline RL, we follow (Setlur et al., 2024; Zhang et al., 2025a; Team et al., 2025) and define pairwise comparisons based on correctness and compactness. Formally, we define: